OpenStudy (anonymous):

solve the system of equations. Enter the solution as an ordered pair , 4x-7y=-23 & 8x+5y=49

OpenStudy (saifoo.khan):

3, 5

OpenStudy (anonymous):

4x - 7y = -23 8x + 5y = 49 8x - 14 y = -46 -8x - 5y = - 49 add everything together -19y = -95 y = 5 4x - 35 = -23 4x = 12 x = 3 (3,5)